which grows faster as x increases, x^3 of 3^x? Explain with examples.
Answer by jim_thompson5910(28536)

3^x grows faster because from x = 4 to x = 5, we get the following

x = 4:
x^3 = 4^3 = 64
3^x = 3^4 = 81

---------------------
x = 5

x^3 = 5^3 = 125
3^x = 3^5 = 243


So as x goes from 4 to 5, x^3 goes from 64 to 125. At the same time, 3^x goes from 81 to 243.

Subtract each result:

For x^3: 125 - 64 = 61

For 3^x: 243 - 125 = 118

and you can see that the change from x = 4 to x = 5 is much larger for 3^x

This shows us that 3^x grows faster as x increases.